# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ 3 Isogonal shear and tissue shear modulus
As discussed in the main text, and shown in Fig. 8 our simulations show that a non-zero cell-level shear modulus is required for extension via active T1s. Here, we analyze the relation of cell-level and tissue-level shear modulus. Crucially, because of the dominance of cortical tensions, a tissue patch in our model will respond to externally applied forces via an isogonal deformation. We first show that an isogonal deformation can create tissue-scale shear (on the level of cell centroid displacement – the transformation of cell vertices is necessarily non-affine to preserve angles). Then we show that these shear modes correspond to the response of the tissue to external force by computing the Hessian of our cell elastic energy in the subspace spanned by isogonal deformations, and measure the shear modulus.
@@ 3.1 Quadratic isogonal mode profile creates shear deformation
To show that isogonal modes can cause shear deformations, not just dilation/contraction of cells, we make use of the isogonal mode parametrization introduced in Ref. [13]. It assigns an isogonal “potential” $\Theta_{i}$ to each cell, and calculates the cell displacements from the $\Theta_{i}$ and the edge tension vectors. In the following, we will show that a constant gradient in the isogonal potential generates a uniform translation in real space. By integration, this implies that a quadratic spatial profile of the isogonal potential creates a shear.
@@
Now we aim to show that a constant gradient in $\Theta_{i}$ drives a uniform displacement of the $\mathbf{r}_{ijk}$. Specifically, by uniform gradient we mean $\Theta_{i}=\mathbf{t}_{i}.\mathbf{a}$, i.e. a linear gradient in the tension space ($\mathbf{t}_{i}$ is the position of the tension triangulation vertex corresponding to cell $i$, such that $\hat{\mathbf{T}}_{ij}=\mathbf{t}_{j}-\mathbf{t}_{i}$). To show that the displacement in real space is uniform, it is enough to show that two adjacent vertices are displaced identically. By induction, this implies that all displacements are identical. It is therefore sufficient to consider a quartet of cells ($i=1{-}4$), corresponding to a “kite” in tension space (note that $\hat{\mathbf{a}}.\mathbf{b}$ is identical to the wedge product $\mathbf{a}\wedge\mathbf{b}$.)
@@
To show that these displacements are identical, we project them onto two conveniently chosen, linearly independent vectors, namely $\hat{\mathbf{T}}_{12}$ and $\hat{\mathbf{T}}_{13}$. For the latter we find
@@
To show equality of these two right-hand sides, we use that given $\hat{\mathbf{T}}_{13}.\mathbf{a}$ and $\hat{\mathbf{T}}_{14}.\mathbf{a}$ we can find $\mathbf{a}$ and substitute the result into $\hat{\mathbf{T}}_{12}.\mathbf{a}$. We start by “expanding the identity”
@@ 3.2 Isogonal Hessian and tissue-scale isogonal shear modulus
Above we have shown that isogonal modes can create a shear (on the level of cell centroids). We now show (a) that shear transformations are the lowest-energy isogonal modes and hence will dominate the response of the tissue to external forces, and (b) that shear transformations incur a finite energy cost, implying a finite shear modulus – the tissue is not a fluid. To do this, we assume that the tissue is initially at the elastic energy minimum under the constraints of a given tension tension triangulation. In this case, we can approximate the energy $E$ by a quadratic function of displacement by computing the Hessian. In our model we impose that vertex angles are given by the intrinsic tension triangulation, hence we cannot study how tensions respond to external forces. Further, due to the dominance of cortical tensions, the lowest-energy deformations of the tissue are isogonal. Hence, here, we only allow isogonal displacements, and we will project the Hessian onto the space of isogonal modes. The lowest eigenvectors of the isogonal Hessian will then correspond to the modes excited by external forcing.
The isogonal unit vectors convert the isogonal potentials into real-space displacement of vertices: the isogonal potential $\Theta_{j}$ at cell $j$ contributes a term $\Theta_{j}\mathbf{I}_{j}^{jkl}$ to the displacement of vertex $\mathbf{r}_{jkl}$ adjacent to $j$. The $\mathbf{I}_{j}^{jkl}$ are defined by the action of isogonal modes Eq. (S7) and serve to project the Hessian onto the space of isogonal modes. Explicitly, $I_{j}^{jkl}=\tfrac{1}{S_{jkl}}\mathbf{T}_{kl}$, where $S_{jkl}=\hat{\mathbf{T}}_{jk}.\mathbf{T}_{kl}$ is the area of the tension triangle $(jkl)$. The isogonal Hessian reads:
@@
Note that only the $E_{\mathcal{C}}$ term from Eq.1 contributes to the isogonal Hessian, by virtue of the isogonal projection.
@@
where $s(h)$ is the shear induced by the vertex displacement. In practice, we compute $s(h)$ by least-squares fit of a shear transformation to the displacement map $\mathbf{r}_{ijk}\mapsto\mathbf{r}_{ijk}+h\cdot\delta\mathbf{r}_{ijk}$). We then fit a parabola to the curve $(s(h),E_{\mathcal{C}}(h))$ to extract the $2^{\mathrm{nd}}$ derivative. We can then investigate how the isogonal shear modulus depends on the cell-level shear modulus in the elastic energy. As expected, we find that the two are proportional (Fig. S13B), showing that the tissue as a whole is solid, in the sense of a non-vanishing shear modulus. We note that a proper investigation of the tissue-scale shear modulus will require addressing the question of how tensions change when external forces are applied, and is hence beyond the scope of the paper.
Figure S13: Cell-level shear modulus determines tissue level shear modulus. A Large-scale tissue shear are the lowest energy isogonal modes. Reference configuration (minimum energy for a given tension triangulation, left) and deformed configuration (right, outline shows reference configuration), deformed according to the isogonal Hessian eigenvector of one of the two lowest eigenvalues in an example tissue patch. B Tissue level “isogonal shear modulus” (see Eq. (S20) for a definition) is proportional to cell level shear modulus. Tissue level isogonal shear modulus vs cell-level shear modulus for the tissue patch shown in A.
@@ 4 Model details and simulation methods
In the following, we explain the simulations presented in the paper and provide the modeling and implementation details. These comprise the mean-field calculation of the dynamics of the distribution of vertex angles, the symmetric lattice of a single intercalating quartet (shown in the companion paper [17]), and the tissue-scale simulation of a disordered epithelial tissue comprising both active and passive regions. All code used to create and analyse the simulations is available on GitHub https://github.com/nikolas-claussen/CE_simulation_public.
@@ 4.1 Positive feedback tension dynamics and single-vertex simulation
We begin by describing the intrinsic tension dynamics of a single tension triangle. This is the basis for the singe-triangle simulation in Fig. 5, where we show the dynamics of the distribution of vertex angles predicted by a positive tension feedback model. Under this model, the edge tensions $\tilde{T}_{\alpha}$, $\alpha=1{-}3$, at a vertex evolve according to:
@@
where $n>1$ and $\tau_{\mathrm{T}}$ is a time scale converting simulation time into minutes. Note that we fix the overall tension scale by the constraint that the triangle perimeter $P=\sum_{\alpha}\tilde{T}_{\alpha}$ is fixed (corresponding to a finite total myosin pool). This simplified the analysis of saturating positive feedback below, and is consistent with the single-quartet simulations in the companion paper. Fixing the triangle area $A$ instead does not significantly affect the results for the single triangle simulations. The feedback exponent was taken as $n=4$ as in Ref. [17]. Different values give very similar results, so no systematic optimization was performed. We note that since the tension constraint $P=\mathrm{const.}$ or $A=\mathrm{const.}$ only affects the overall triangle size, and the trajectories in LTC space (but not their speed) are independent of it.
@@
To find the trajectories in the LTC space, we integrated Eq. (5). A change of variables from side length to the LTC space yields the flow shown in Fig. 5A. To model the effect of intercalations, which remove highly obtuse tension triangles from the distribution, we stopped integration once the maximum relative tension increased over $T_{\mathrm{rel}}=1.56$, the median of the relative tension at the moment of a T1 pbserved the tissue-scale simulations with isogonal stretching (as well as in the experimental data). These “collapse” triangles were removed from the simulation ensemble.
@@
To determine the tension timescale $\tau_{T}$, we compute the dynamics of the relative tension (the ratio of the edge tension to that of the adjacent edges) for all edges that undergo a T1 transition and align them temporally based on the time the T1 process occurs. This leads an average $T_{\mathrm{rel}}(t)$ that can be compared to the time traces from the experimental data reported in the companion paper[17] to fit $\tau_{T}$. Because the tension dynamics in the tissue-scale simulations is affected by tension balancing (see below), while the single-triangle and single-quartet simulations are not, we fit $\tau_{T}$ separately for the two classes of simulations. Note that we fit the tension timescale to a microscopic process, the T1 transition, and the resulting time course for macroscopic observables (tissue extension, LTC distribution) matches the experimental data without additional adjustment.
@@ 4.2 Tension dynamics post intercalation: Myosin handover and passive tension
As explained in the companion paper, we propose a myosin handover mechanism to explain the extension of the new interface post intercalation. A interface with cortical tension $T$ is comprised of the adherens-junctional actomyosin cortex of the two adjacent cells, which are coupled mechanically via adherens junctions. Under force balance, the total tension $T$ has to be constant along the cortex, but the individual tensions on either side can be non-uniform, as the resulting traction forces are exchanged via adherens junctions. In the following, we assume as a first order approximation that the level of active tension (i.e. myosin concentration) varies linearly along an interface (similar calculations have been performed in Ref. [73] to calculate interfacial shear stress). This allows us to geometrically obtain the myosin concentration at the individual cortices that will form the two juxtaposed sides of the new interface (see companion paper).
Consider a vertex, where interfaces $i=0,1,2$ (with tensions $T_{0},T_{1},T_{2}$) meet, and let 00 be the interface about to collapse. For an illustration, see Ref. [17], Fig. 4). The three cells that meet at the vertex will be referred to by $(01),(12),(21)$ (where e.g. $(01)$ is the cell abutting interfaces 00 and $1$). Let $m_{01},m_{12},m_{21}$ be the motor molecule concentrations (in units of tension) at the vertex in the junctional cortices of the three cells. Then, the tensions are related to the motor molecule concentrations as:
@@
This uses the assumption of myosin continuity at vertices, and the fact that the tension on an interface is the sum of the tensions of the two cortices that make it up. The motor molecule concentration on the cortex belonging to the new interfaces, post collapse, will be equal to $m_{12}$. Solving for this in terms of the tensions:
@@
The new interface consists of two cortices, coming from the two vertices of the collapsed interface. Let the tensions at the two triangles be $T_{0},T_{1},T_{2}$ and $T_{0},T_{1}^{\prime},T_{2}^{\prime}$. Let $T_{a}$ be the active tension on the new interface immediately after the T1. It is the equal to
@@
Note however that the total tension $T_{n}$ on the new interface is not necessary equal to $T_{a}$. The total tension is defined geometrically from the angles at the new interface (or, equivalently, the tension triangulation vertices). Indeed, generally, $T_{n}>T_{a}$, i.e. the active tension on the new interface is not enough to balance the tension due to the adjacent edges. As explained in the companion paper, we introduce a passive tension $T_{p}$ on the new edge which balances this deficit
@@
For example, if a perfectly symmetric quartet collapses when the vertex angle facing the collapsing edge is $90^{\circ}$, $T_{1}=T_{2}=T_{1}^{\prime}=T_{2}^{\prime}=1$ and $T_{0}=T_{n}=\sqrt{2}$. Therefore, $T_{a,n}=2-\sqrt{2}\approx 0.6$ and $T_{p,n}=\sqrt{2}-(2-\sqrt{2})\approx 0.8$. Note that by the triangle inequality, for any convex quadrilateral with perimeter $P$ and diagonals $D_{1},D_{2}$, one has $P/2\leq D_{1}+D_{2}\leq P$. Applying this to the quadrilateral formed by the two tension triangles at the collapsing interface, we get $T_{a,n}\geq 0$ and $T_{p,n}\geq 0$: the handover formula always results in positive active and passive tensions. Further, the “handover” mechanism robustly generates irreversible T1s: if a interface were to collapse back after a T1, the newly formed interface would inherit high myosin levels and therefore be likely to collapse again.
@@
Figure S14: Absence of energy barrier to intercalations A Length of the central interface in a quartet for a prescribed isosceles tension triangle parametrized by the angle $\beta$ (the isosceles property implies $\alpha=(\pi-\beta)/2$). B The solid line shows the interface length in the Voronoi tessellation $\ell_{\mathrm{ref}}$ (cf. Eq. (10)). The interface length obtained by minimizing elastic energy Eq. (3) with isotropic target shape (dot-dashed line) closely follows the Voronoi length and vanishes at the same critical angle. By contrast, minimizing the “area–perimeter” energy of the vertex-model $E_{\mathrm{cell}}\sim(A-A_{0})^{2}+(P-P_{0})^{2}$ in the solid phase, $P_{0}/\sqrt{A_{0}}<3.81$, gives a interface length (dashed black line) that vanishes only for $\beta\to\pi$ and the energy diverges in this limit (dashed green line).

### v3

@@ 4 Isogonal shear and tissue shear modulus
As discussed in the main text, and shown in Fig. 7 our simulations show that a non-zero cell-level shear modulus is required for extension via active T1s when the tissue extends against external restoring force (e.g. from adjacent tissues). Here, we analyze the relation between cell-level and tissue-level shear modulus. Crucially, because of the dominance of cortical tensions, a tissue patch in our model will respond to externally applied forces via an isogonal deformation. We first show that an isogonal deformation can create tissue-scale shear (on the level of cell centroid displacement – the transformation of cell vertices is necessarily non-affine to preserve angles). Then we show that these shear modes correspond to the response of the tissue to external force by computing the Hessian of our cell elastic energy in the subspace spanned by isogonal deformations, and measure the shear modulus. Our discussion here applies to deformations of the bulk of a tissue; special care is needed when considering deformations that change the tissue boundary, as we explain below.
@@ 4.1 Quadratic isogonal mode profile creates shear deformation
To show that the isogonal modes can cause shear deformations, not just dilation/contraction of cells, we make use of the isogonal mode parametrization introduced in Ref. [Noll.etal2017]. It assigns an isogonal “potential” $\Theta_{i}$ to each cell, and calculates the cell displacements from the $\Theta_{i}$ and the edge tension vectors. In the following, we will show that a constant gradient in the isogonal potential generates a uniform translation in real space. By integration, this implies that a quadratic spatial profile of the isogonal potential creates a shear.
@@
Now we aim to show that a constant gradient in $\Theta_{i}$ drives a uniform displacement of the $\mathbf{r}_{ijk}$. Specifically, by uniform gradient, we mean $\Theta_{i}=\mathbf{t}_{i}.\mathbf{a}$, i.e. a linear gradient in the tension space ($\mathbf{t}_{i}$ is the position of the tension triangulation vertex corresponding to cell $i$, such that $\hat{\mathbf{T}}_{ij}=\mathbf{t}_{j}-\mathbf{t}_{i}$). To show that the displacement in real space is uniform, it is enough to show that two adjacent vertices are displaced identically. By induction, this implies that all displacements are identical. It is therefore sufficient to consider a quartet of cells ($i=1{-}4$), corresponding to a “kite” in tension space (note that $\hat{\mathbf{a}}.\mathbf{b}$ is identical to the wedge product $\mathbf{a}\wedge\mathbf{b}$.)
@@
To show that these displacements are identical, we project them onto two conveniently chosen, linearly independent vectors, namely $\hat{\mathbf{T}}_{12}$ and $\hat{\mathbf{T}}_{13}$. For the latter, we find
@@
To show the equality of these two right-hand sides, we use that given $\hat{\mathbf{T}}_{13}.\mathbf{a}$ and $\hat{\mathbf{T}}_{14}.\mathbf{a}$ we can find $\mathbf{a}$ and substitute the result into $\hat{\mathbf{T}}_{12}.\mathbf{a}$. We start by “expanding the identity”
@@ 4.2 Isogonal Hessian and tissue-scale isogonal shear modulus
Above we have shown that isogonal modes can create a shear (on the level of cell centroids). We now show (a) that shear transformations are the lowest-energy isogonal modes and hence will dominate the response of the tissue to external forces, and (b) that shear transformations incur a finite energy cost, implying a finite shear modulus – the tissue is not a fluid. To do this, we assume that the tissue is initially at the elastic energy minimum under the constraints of a given tension tension triangulation. In this case, we can approximate the energy $E$ by a quadratic function of displacement by computing the Hessian. In our model we impose that vertex angles are given by the intrinsic tension triangulation, hence we cannot study how tensions respond to external forces. Further, due to the dominance of cortical tensions, the lowest-energy deformations of the tissue are isogonal. Hence, here, we only allow isogonal displacements, and we will project the Hessian onto the space of isogonal modes. The lowest eigenvectors of the isogonal Hessian will then correspond to the modes excited by external forcing. Note that isogonal deformations only leave the dominant junctional elastic energy $\sum_{ij}T_{ij}\ell_{ij}-p\sum_{i}A_{i}$ invariant if they don’t deform the tissue boundary. The tissue patch (shown in Fig. S13A) we consider in the following calculation should therefore be imagined as embedded into a larger cell array with overall fixed boundaries, as in the simulations of Fig. 7 where the net shear of active and passive regions is zero.
The isogonal unit vectors convert the isogonal potentials into real-space displacement of vertices: the isogonal potential $\Theta_{j}$ at cell $j$ contributes a term $\Theta_{j}\mathbf{I}_{j}^{jkl}$ to the displacement of vertex $\mathbf{r}_{jkl}$ adjacent to $j$. The $\mathbf{I}_{j}^{jkl}$ are defined by the action of isogonal modes (S8) and serve to project the Hessian onto the space of isogonal modes. Explicitly, $I_{j}^{jkl}=\tfrac{1}{S_{jkl}}\mathbf{T}_{kl}$, where $S_{jkl}=\hat{\mathbf{T}}_{jk}.\mathbf{T}_{kl}$ is the area of the tension triangle $(jkl)$. The isogonal Hessian reads:
@@
Note that only the $E_{\mathcal{C}}$ term from (1) contributes to the isogonal Hessian, by virtue of the isogonal projection.
@@
where $s(h)$ is the shear induced by the vertex displacement. In practice, we compute $s(h)$ by least-squares fit of a shear transformation to the displacement map $\mathbf{r}_{ijk}\mapsto\mathbf{r}_{ijk}+h\cdot\delta\mathbf{r}_{ijk}$). We then fit a parabola to the curve $(s(h),E_{\mathcal{C}}(h))$ to extract the $2^{\mathrm{nd}}$ derivative. We can then investigate how the isogonal shear modulus depends on the cell-level shear modulus in the elastic energy. As expected, we find that the two are proportional (Fig. S13B), showing that the tissue as a whole is solid, in the sense of a non-vanishing shear modulus. We note that a complete investigation of the tissue-scale shear modulus, i.e. a complete account of tissue rheology, will require addressing (a) energy costs due to deformations of the tissue boundary and (b) the question of how tensions change when external forces are applied and is hence beyond the scope of the paper.
@@ 5 Model details and simulation methods
In the following, we explain the simulations presented in the paper and provide the modeling and implementation details. These comprise the mean-field calculation of the dynamics of the distribution of vertex angles, the symmetric lattice of a single intercalating quartet (shown in the companion paper [Brauns.etal2024elife]), and the tissue-scale simulation of a disordered epithelial tissue comprising both active and passive regions. All code used to run and analyze the simulations is available on GitHub https://github.com/nikolas-claussen/CE_simulation_public.
@@ 5.1 Positive feedback tension dynamics and single-vertex simulation
We begin by describing the intrinsic tension dynamics of a single tension triangle. This is the basis for the singe-triangle simulation in Figs. 4B and 4C, where we show the dynamics of the distribution of vertex angles predicted by a positive tension feedback model. Under this model, the edge tensions $\tilde{T}_{\alpha}$, $\alpha=1{-}3$, at a vertex evolve according to:
@@
where $n>1$ and $\tau_{\mathrm{T}}$ is a time scale converting simulation time into minutes. Note that we fix the overall tension scale by the constraint that the triangle perimeter $P=\sum_{\alpha}\tilde{T}_{\alpha}$ is fixed (corresponding to a finite total myosin pool). This simplified the analysis of saturating positive feedback below and is consistent with the single-quartet simulations in the companion paper. Fixing the triangle area $A$ instead does not significantly affect the results for the single triangle simulations. The feedback exponent was taken as $n=4$ as in Ref. [Brauns.etal2024elife]. Different values give very similar results, so no systematic optimization was performed. We note that since the tension constraint $P=\mathrm{const.}$ or $A=\mathrm{const.}$ only affects the overall triangle size, the trajectories in LTC space (but not their speed) are independent of it.
@@
To find the trajectories in the LTC space, we integrated (5). A change of variables from side length to the LTC space yields the flow shown in Fig. 4C. To model the effect of intercalations, which remove highly obtuse tension triangles from the distribution, we stopped integration once the maximum relative tension increased over $T_{\mathrm{rel}}=1.56$, the median of the relative tension at the moment of a T1 observed the tissue-scale simulations with isogonal stretching (as well as in the experimental data). These “collapse” triangles were removed from the simulation ensemble.
@@
To determine the tension timescale $\tau_{T}$, we compute the dynamics of the relative tension (the ratio of the edge tension to that of the adjacent edges) for all edges that undergo a T1 transition and align them temporally based on the time the T1 process occurs. This leads an average $T_{\mathrm{rel}}(t)$ that can be compared to the time traces from the experimental data reported in the companion paper[Brauns.etal2024elife] to fit $\tau_{T}$. Because the tension dynamics in the tissue-scale simulations are affected by tension balancing (see below), while the single-triangle and single-quartet simulations are not, we fit $\tau_{T}$ separately for the two classes of simulations. Note that we fit the tension timescale to a microscopic process, the T1 transition, and the resulting time course for macroscopic observables (tissue extension, LTC distribution) matches the experimental data without additional adjustment.
@@ 5.2 Tension dynamics post intercalation: Myosin handover and passive tension
As explained in the companion paper, we propose a myosin handover mechanism to explain the extension of the new interface post-intercalation. An interface with cortical tension $T$ is comprised of the adherens-junctional actomyosin cortex of the two adjacent cells, which are coupled mechanically via adherens junctions. Under force balance, the total tension $T$ has to be constant along the cortex, but the individual tensions on either side can be non-uniform, as the resulting traction forces are exchanged via adherens junctions. In the following, we assume as a first-order approximation that the level of active tension (i.e. myosin concentration) varies linearly along an interface (similar calculations have been performed in Ref. [Kale.etal2018] to calculate interfacial shear stress). This allows us to geometrically obtain the myosin concentration at the individual cortices that will form the two juxtaposed sides of the new interface (see companion paper).
Consider a vertex, where interfaces $i=0,1,2$ (with tensions $T_{0},T_{1},T_{2}$) meet, and let 00 be the interface about to collapse. For an illustration, see Ref. [Brauns.etal2024elife], Fig. 4). The three cells that meet at the vertex will be referred to by $(01),(12),(21)$ (where e.g. $(01)$ is the cell abutting interfaces 00 and $1$). Let $m_{01},m_{12},m_{21}$ be the motor molecule concentrations (in units of tension) at the vertex in the junctional cortices of the three cells. Then, the tensions are related to the motor molecule concentrations as
@@
This uses the assumption of myosin continuity at vertices and the fact that the tension on an interface is the sum of the tensions of the two cortices that make it up. The motor molecule concentration on the cortex belonging to the new interfaces post-collapse will be equal to $m_{12}$. Solving for this in terms of the tensions:
@@
The new interface consists of two cortices, coming from the two vertices of the collapsed interface. Let the tensions at the two triangles be $T_{0},T_{1},T_{2}$ and $T_{0},T_{1}^{\prime},T_{2}^{\prime}$. Let $T_{a}$ be the active tension on the new interface immediately after the T1. It is equal to
@@
Note however that the total tension $T_{n}$ on the new interface is not necessarily equal to $T_{a}$. The total tension is defined geometrically from the angles at the new interface (or, equivalently, the tension triangulation vertices). Indeed, generally, $T_{n}>T_{a}$, i.e. the active tension on the new interface is not enough to balance the tension due to the adjacent edges. As explained in the companion paper, we introduce a passive tension $T_{p}$ on the new edge which balances this deficit
@@
For example, if a perfectly symmetric quartet collapses when the vertex angle facing the collapsing edge is $90^{\circ}$, $T_{1}=T_{2}=T_{1}^{\prime}=T_{2}^{\prime}=1$ and $T_{0}=T_{n}=\sqrt{2}$. Therefore, $T_{a,n}=2-\sqrt{2}\approx 0.6$ and $T_{p,n}=\sqrt{2}-(2-\sqrt{2})\approx 0.8$. Note that by the triangle inequality, for any convex quadrilateral with perimeter $P$ and diagonals $D_{1},D_{2}$, one has $P/2\leq D_{1}+D_{2}\leq P$. Applying this to the quadrilateral formed by the two tension triangles at the collapsing interface, we get $T_{a,n}\geq 0$ and $T_{p,n}\geq 0$: the handover formula always results in positive active and passive tensions. Further, the “handover” mechanism robustly generates irreversible T1s: if an interface were to collapse back after a T1, the newly formed interface would inherit high myosin levels and therefore be likely to collapse again.